Presented below is information related to Bramble Manufacturing Corporation.

Asset

Cost

Estimated Salvage

Estimated Life (in years)

A

$47,900 $5,500 10

B

33,400 4,600 9

C

36,300 3,900 9

D

18,400 2,300 7

E

24,600 2,400 6

Compute the rate of depreciation per year to be applied to the plant assets under the composite method. (Round answer to 2 decimal place, e.g. 4.83%.)

Composite rate

enter the composite rate in percentages rounded to 2 decimal places

%

Prepare the adjusting entry necessary at the end of the year to record depreciation for the year. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Use Plant Assets related account.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

Prepare the entry to record the sale of asset D for cash of $4,500. It was used for 6 years, and depreciation was entered under the composite method. (If no entry is required, select "No entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Use Plant Assets related account.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount
Answer #1

Composite method of depreciation is a kind of group depreciation. Composite depreciation will base calculations for all asset categories in one group in straight line method. The entity bramble manufacturing corporation has to calculate composite rate of depreciation and composite life.

Assuming the A,B,C,D,E all belong to same category of assets (e.g. plant & machiney or Furniture & Fittings) the composite rate and composite life is calculated as below.

(Amount in $)
Asset Cost
(A)
Estimated Salvage value
(B)
Depreciable Amount
C=(A-B)
Depreciable life
(D)
Straight line depreciation per annum
A 47900 5500 42400 10 4240
B 33400 4600 28800 9 3200
C 36300 3900 32400 9 3600
D 18400 2300 16100 7 2300
E 24600 2400 22200 6 3700
Total 160600 18700 141900 17040

Composite rate of depreciation to be applied on plant assets= Total Depreciation per annum /Total Original cost

=17040/160600 = 10.61%

Composite life = Total depreciable amount/ depreciation per year

=141900/17040 = 8.33 Years rounded to 8 years

Composite depreciation expense each year equals the composite depreciation rate applied to the total original cost.Composite depreciation for each of five reporting years is(Before asset D is sold) :

Composite depreciation = (0.1061)*160600= 17040.

Accounting entry at end of the year to record depreciation is as follows:-

S.no Account Title Debit Credit
1 Depreciation expense Account( on plant assets) 17040
Accumulated depreciation Account 17040

Composite method do not allow any recognition of profit or loss on sale of assets.So when an entity sells assets

  • firstly debit the cash received.
  • Secondly credit the asset account (original cost)
  • Any difference to be routed through accumulated depreciation account.

Hence in the given case Asset D is sold for $4,500 , the following entry to be passed in books of accounts.

S.no Account Title Debit Credit
1 Bank account 4500
Accumulated depreciation Account 13900
Asset D account 18400